OpenStudy (anonymous):

Where is the missing $1? You saw a shirt for 97 dollars. You borrowed 50 from your mom and 50 from dad. After you bought the shirt you were leftt with 3 dollars change. You give each parent a dollar and kept the other. Now you owe each parent 49. (49) (2) + 1= 99. Where is the missing 1 dollar?

OpenStudy (sasogeek):

Ok so if you're giving each parent their 50 dollars back, then that would be (49)(2) + 2 = 100 +2 because you already gave them 1 dollar each already. each means 1 + 1. check this out 49 + 1 = 50 father gets his 50 dollars back 49 + 1 = 50 mother gets her 50 dollars back you have 1 extra dollar with you. so in essence, you logic wasn't well worked out, especially with the last equation ;)

OpenStudy (sasogeek):

All things being equal, if we should go by the equation you gave, then the extra dollar is with you